### v1

# Mapping Trustworthiness in Large Language Models: A Bibliometric Analysis Bridging Theory to Practice

Abstract.
The rapid proliferation of Large Language Models (LLMs) has raised pressing concerns regarding their trustworthiness, spanning issues of reliability, transparency, fairness, and ethical alignment. Despite the increasing adoption of LLMs across various domains, there remains a lack of consensus on how to operationalize trustworthiness in practice. This study bridges the gap between theoretical discussions and implementation by conducting a bibliometric mapping analysis of 2,006 publications from 2019 to 2025. Through co-authorship networks, keyword co-occurrence analysis, and thematic evolution tracking, we identify key research trends, influential authors, and prevailing definitions of LLM trustworthiness. Additionally, a systematic review of 68 core papers is conducted to examine conceptualizations of trust and their practical implications. Our findings reveal that trustworthiness in LLMs is often framed through existing organizational trust frameworks, emphasizing dimensions such as ability, benevolence, and integrity. However, a significant gap exists in translating these principles into concrete development strategies. To address this, we propose a structured mapping of 20 trust-enhancing techniques across the LLM lifecycle, including retrieval-augmented generation (RAG), explainability techniques, and post-training audits. By synthesizing bibliometric insights with practical strategies, this study contributes towards fostering more transparent, accountable, and ethically aligned LLMs, ensuring their responsible deployment in real-world applications.
Trustworthiness, AI ethics, Large Language Models, Bibliometric Mapping
copyright: acmlicensedjournalyear: 1014doi: XXXXXXX.XXXXXXXconference: The Xth International Conference on XXXX; XX–$X$ June, 2025; XXXX, XXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: General and reference Surveys and overviewsccs: Software and its engineering Software development techniquesccs: Social and professional topics Computing / technology policy
1. Introduction
Artificial Intelligence (AI) has become a cornerstone of technological progress, yet remains at the centre of ethical debate. More than 200 guidelines and principles have been proposed to guide its development and use, reflecting the growing recognition of its societal impact (Hagendorff, 2020; Corrêa et al., 2023). These AI ethical guidelines, which have been proposed by public organisations, society, academia and private companies, do not provide objective guidance to developers, but rather abstract principles that are not practical for developers to operationalise, e.g, transparency, privacy, justice & fairness (Vakkuri et al., 2021; Floridi and Cowls, 2022; de Cerqueira et al., 2022).
As an evolution of AI system that employs advanced deep learning techniques, Large Language Models (LLMs) are trained on vast datasets and capable of generating novel text, images, audio, and video, based on the data it was trained on (Sun et al., 2024). Their rapid adoption has sparked both enthusiasm and concern, as their potential for negative outcomes – such as spreading misinformation, perpetuating biases, or hallucination – has become evident (Liu et al., 2023; Wang et al., 2023). Despite efforts like the Paris AI Action Summit (Élysée, 2025), which seeks to promote responsible AI development and international cooperation, significant challenges remain. For instance, Google’s recent decision to remove its AI principles prohibiting the use of AI in autonomous weapons and surveillance has raised questions about the commitment of major tech companies to ethical AI (Hooker and Vallance, 2025). This shift underscores the tension between ethical ideals and commercial or geopolitical interests, highlighting the need for robust, enforceable legislations.
The swift popularization of LLMs have encouraged researchers in understanding the specific ethical dimensions of this new generative AI (genAI), introducing them mainly as trustworthiness taxonomies, i.e., a rebranded set of similar abstract AI ethical principles (Sun et al., 2024; Liu et al., 2023). Trustworthiness in LLMs can be characterized in terms of reliability, safety, fairness, resistance to misuse, explainability & reasoning, social norm and robustness (Liu et al., 2023). Thus, the focus of AI ethics discussions has shifted to the trustworthiness of LLM systems. The current debate on LLM trustworthiness largely focus on providing definitions, benchmarks and evaluation methods (Sun et al., 2024; Liu et al., 2023; Wang et al., 2023).
Analogously with AI ethics, the proliferation of trustworthiness taxonomies for LLMs has not led to a clear consensus on definitions or a practical framework to guide the development of these systems (Smith et al., 2025). Despite the growing body of literature on LLM trustworthiness, there is still no unified set of practical guidelines to help developers mitigate risks throughout the development and deployment process. This gap between theory and practice not only hinders the creation of more trustworthy LLM systems but also masks deeper structural problems that have persisted since the early days of AI ethics.
This study aims to bridge the gap between theory and practice by investigating the means of operationalising trustworthiness in LLMs. We conduct a bibliometric mapping study, with 2,006 studies from 2019 to 2025, to identify the most influential authors, research trends and approaches and systematically analyse their similarities and differences. Additionally, by manually reading 68 key documents, we also conduct an analysis of the different definitions of trustworthiness proposed in the literature, providing insights into how different authors conceptualise trust in LLMs. Furthermore, we extend our analysis by identifying practical strategies that developers can use to enhance the trustworthiness of LLMs throughout the development lifecycle – from pre-training to deployment and real-world use. By bringing these perspectives together, we aim to provide both a theoretical and practical basis for promoting more ethical and trustworthy LLMs.
In order to guide this study, we devised the following Research Questions:
  • RQ1: What are the research trends, key authors, and thematic areas in the study of LLM trustworthiness?
  • RQ2: What are the LLM trust/trustworthiness definitions proposed by different authors in the literature?
  • RQ3: What practical techniques and tools can be used to enhance LLM trustworthiness throughout the LLM lifecycle?
The main contributions of this study are as follows:
  • Bibliometric Mapping Analysis: Using Bibliometrix, an R-tool, we conduct a bibliometric mapping study to identify the most influential authors, definitions, and approaches in the field. This analysis reveals trends, patterns, and areas of divergence in the literature on LLM trustworthiness.
  • Systematic Analysis of Trustworthiness Definitions: We compile and categorize various definitions of trust and trustworthiness in LLMs proposed by different authors. By analyzing these definitions, we identify commonalities, differences, and conceptual overlaps with related terms such as Responsible AI and AI Ethics.
  • LLM Trustworthiness Enhancing Strategies: We propose a mapping of practical techniques, technologies and processess that developers can use to enhance LLM trustworthiness at various stages of the development lifecycle, from pre-training to post-training, deployment and end-use. This helps in bridging the gap between abstract taxonomies and actionable steps.
  • Actionable Recommendations: Based on our findings, we offer actionable recommendations for developers, policymakers, and researchers to operationalize trustworthiness principles in practice. These recommendations aim to address the structural challenges that have hindered the development of trustworthy AI systems.
By integrating a systematic review of trustworthiness definitions with bibliometric insights and practical techniques, this work contributes to ongoing efforts to operationalize trustworthiness in LLMs and offers concrete steps toward the development of more trustworthy LLM systems.
This article is organized as follows: in the Section 2, we review similar works that conduct literature reviews on AI ethical guidelines and LLM trustworthiness taxonomies, highlighting their similarities and differences. In Section 3, we describe the Bibliometrix methodology adopted. In Section 4, we present the main findings, including systematic analyses of LLM trustworthiness definitions and trustworthiness-enhancing techniques. In Section LABEL:threats_to_validity, we discuss the study’s limitations and potential biases. Finally, in Section LABEL:final_remarks_section, we reflect on the practical and theoretical implications of the results, suggesting future directions for research and development.
2. Related Work
In the context of this study, we examine previous literature reviews related to LLM and genAI trust/trustworthiness and ethics in the context of this study. Among the most relevant works is that of Liu et al. (Liu et al., 2023), which proposes a taxonomy for evaluating trustworthiness in LLM in seven major categeories: reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness. Although the paper is not a literature review, it does describe each of the major categories with examples. However, the authors fail to provide a clear definition of what is trustworthy and do not provide practical guidance for developers on how to operationalise their taxonomy.
Albahri et al. (Albahri et al., 2023) conducted a systematic science mapping analysis in the context of trustworthy and explainable AI (XAI) in healthcare systems, and found various XAI techniques for developers to operationalise. Unlike their approach, we aim to provide means to operationalise more of the trustworthiness categories.
It can be seen that many scholars have provided reviews of AI ethical principles, such as (Mittelstadt, 2019), (Jobin et al., 2019), (Corrêa et al., 2023), (Hagendorff, 2020), (Siqueira De Cerqueira et al., 2021), also providing an overview of publicly available tools that can help operationalise ethical principles in AI (Morley et al., 2019a). However, they are not as relevant to our review as they appeared before the LLM and genAI hype, and we want to explore the notion of trustworthiness in LLM and genAI and how it relates to ethics in AI, also providing practical guidance for developers to operationalise it.
3. Methodology
This study employs a bibliometric analysis to systematically investigate the landscape of research on trustworthiness in LLMs. Bibliometric analysis is a quantitative research method used to identify trends, influential works, and key research themes in a given domain (Donthu et al., 2021). To achieve this, we utilize Bibliometrix, an open-source R package specifically designed for bibliometric and scientometric studies (Aria and Cuccurullo, 2017).
3.1. Data Collection
The dataset for this study was extracted from the Web of Science (WoS) Core Collection during 21st of February 2025, as this database contain high-quality, peer-reviewed publications in Artificial Intelligence, ethics, and Software Engineering, and provides the highest quality of publications and reliability in its indexing of high-ranking journals. (Caputo et al., 2021). The search query was formulated using a combination of relevant keywords to ensure a comprehensive dataset:
TS=((”Large Language Model*” OR LLM OR ”Generative AI” OR ”genAI”) AND (Trust* OR Ethic*))
Where TS stands for Topic, which encompasses the following fields within a record: Title, Abstract, Author Keywords and Keywords Plus®. The latter is automatically generated by the database, in which its data are words or phrases that frequently appear in the titles of an article’s references, but do not appear in the title of the article itself  [^1] https://support.clarivate.com/ScientificandAcademicResearch/s/article/KeyWords-Plus-generation-creation-and-changes.
The inclusion criteria for the retrieved publications were:
  • Peer-reviewed journal articles and conference proceedings.
  • Papers published in the last 6 years (2019-2025).
  • Studies encompassing discussion on trustworthiness or ethics of LLMs or generative AI.
  • Articles written in English.
After filtering out duplicates, irrelevant papers, and LLM as authors abbreviation, a final dataset of 2,006 publications was obtained. This amount of studies discussing LLMs is impressive, since we had to reduce the year range of our search to only 6 years due to the fact that LLM is a new technology.
Further manual reading: After exploring the broader set with Bibliometrix, we conducted a manual reading of a reduced dataset in order to gain insightful information and cross-validate our findings. This manual reading was undertaken to provide answers to research questions 2 and 3. This minimal set of studies was obtained by filtering within Bibliometrix. An objective way to define whether or not there was a focus on trustworthiness was to look at the discussion at the abstract level. Thus, by filtering papers that had ”trust*” in the abstract, a set of 68 papers was found for manual reading.
3.2. Data Processing and Preprocessing
The raw bibliographic data was exported in BibTeX format and processed using Bibliometrix in R. The following preprocessing steps were conducted:
  • Data Cleaning: Removal of missing entries, duplicate records, and irrelevant fields.
  • Normalization: Unification of author names and keyword variations to ensure consistency.
3.3. Bibliometric Analysis
The bibliometric analysis consisted of the following techniques:
  • Descriptive Analysis: Calculation of publication trends, citation statistics, and author productivity over time.
  • Co-Authorship Network Analysis: Mapping collaborations between authors and institutions to identify key contributors in LLM trustworthiness research.
  • Co-Citation and Bibliographic Coupling: Identifying clusters of highly interconnected research papers and their thematic evolution.
  • Keyword Co-Occurrence Analysis: Investigating the most frequently occurring terms to identify emerging topics and research gaps.
  • Thematic Mapping: Using Biblioshiny (a visualization tool within Bibliometrix) to classify research themes based on centrality and density.
3.4. Visualization and Interpretation
To facilitate a deeper understanding of the research trends, we generated various network visualizations, including:
  • Co-Authorship Networks (to highlight key collaborations).
  • Keyword Co-Occurrence Networks (to reveal thematic clusters).
  • Citation Networks (to track the evolution of influential studies).
The results of these analyses were synthesized to provide insights into the state-of-the-art, research gaps, and future directions in LLM trustworthiness research.
4. Results and Discussion
This study examines the landscape of trustworthiness and ethics research in LLMs and generative AI from 2019 to 2025. A total of 2006 relevant studies, published across 1160 different venues over the past six years, were analyzed. These publications were authored by 7901 researchers, with an average citation count of 8.901 per document. Notably, collaborative research dominates the field, with 96.6% of the studies involving multiple authors (7,632 in total), while only 3.4% were produced by single authors (269 individuals).
4.1. What are the research trends, key authors, and thematic areas in the study of LLM trustworthiness?
4.2. Evolution of Publications from 2019-2025
The study shows a sharp rise in publications on LLM trustworthiness and ethics from 2019 to early 2025, growing from 3 articles in 2019 to 1,434 in 2024, shown in Table 1 and in Figure 1. This growth was driven by advances like GPT-3/4, regulatory debates (e.g., EU AI Act), and societal concerns about AI ethics. Collaborative research dominated, with 96.6% of studies involving multiple authors. By February 2025, 213 articles were published, suggesting continued momentum ( 1,278 projected by year-end). This highlights the increase interest researchers have in trust/trustworthiness and ethical LLM and genAI research.
Table 1. Number of Articles Published Per Year
Year Articles
2019 3
2020 1
2021 4
2022 12
2023 339
2024 1434
2025 213
Figure 1. Evolution of Documents Per Year - 2019 - 2025.
4.3. Most Global Cited Documents
Figure 2 shows the top documents ranked by total number of citations, without any filtering. Articles with high global citations indicate significant impact and influence across disciplines (Donthu et al., 2021).
The document ”ChatGPT for good? On opportunities and challenges of large language models for education” by Kasneci et al. (Kasneci et al., 2023) has received 1,135 global citations. The paper explores the opportunities and challenges of using LLMs like ChatGPT in education, highlighting their potential to personalize learning, create educational content, and support both students and teachers. However, it also addresses critical risks such as bias, over-reliance, ethical concerns, and sustainability, emphasizing the need for responsible integration, teacher training, and ongoing research. The authors advocate for a balanced approach that leverages LLMs to enhance education while ensuring ethical use, transparency, and equitable access. Ongoing efforts to perceive the use of LLMs in education highlights how academia is concerned about this disruptive tool in educational settings.
The second most cited document is ”Opinion Paper: “So what if ChatGPT wrote it?” Multidisciplinary perspectives on opportunities, challenges and implications of generative conversational AI for research, practice and policy” by Dwivedi et al. (Dwivedi et al., 2023), with an equal amount of 1,135 global citations. The paper explores the multidisciplinary impact of ChatGPT, highlighting its potential to enhance productivity across industries while raising concerns about ethics, misinformation, and transparency. It discusses its role in research, business, and education, emphasizing both opportunities and challenges, including AI’s effect on academic integrity and hybrid work. The authors propose multiple future research directions, e.g., on AI governance, human-AI collaboration, and responsible implementation.
The third most cited document ”ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns” by Sallam Malik (Sallam, 2023). This study systematically reviews the potential applications and limitations of ChatGPT in healthcare. It highlights the benefits of ChatGPT in scientific writing, research efficiency, medical education, and clinical workflows, while addressing concerns such as misinformation, bias, ethical and legal challenges, and safety risks. The study emphasises the need for responsible use of AI in healthcare and calls for regulatory guidelines and stakeholder collaboration to ensure ethical and effective implementation.
The three highly cited documents analysed present investigations of the specific adoption and impact of ChatGPT in different domains – education (Kasneci et al., 2023), research, business, and education (Dwivedi et al., 2023), and healthcare (Sallam, 2023). This shows how the research community has responded to the rapid emergence of ChatGPT in several domains. Furthermore, all three top documents have in common the need to ensure the ethical use, transparency and responsible implementation and use of AI.
Figure 2. Most Globally Cited Documents (2022-2024).
4.4. Most Local Cited Sources
When analyzing the most cited sources in our dataset in Table 2, we found an overwhelming 10,237 citations to arXiv preprint documents. This highlights a strong reliance on this preprint repository, suggesting that research in this area is rapidly evolving, with many studies referencing early-stage, non-peer-reviewed work. The prominence of arXiv, a widely used open-access platform for sharing preprints in fields like computer science, physics, and artificial intelligence, reflects the fast-paced nature of this research domain, where immediate access to emerging findings is valuable. However, while arXiV is interesting for the open-access, the documents inside are mostly pre-prints, that is, they mostly lack the scientific rigour that draws from the peer-review process.
In second place is ADV NEUR IN with 1,021 citations, which stands for Advances in Neural Information Processing Systems (NeurIPS). This is one of the most prestigious conferences in artificial intelligence and machine learning, attracting top researchers from around the world. It serves as a key platform for presenting groundbreaking advances in deep learning, reinforcement learning, and computational neuroscience that are shaping the future of AI research. The high citation rate of NeurIPS papers reflects its significant impact, as cutting-edge developments presented at the conference often drive innovation across academia and industry.
Nature comes in third with 980 citations. Nature is one of the world’s most prestigious and influential scientific journals, publishing ground-breaking research in a wide range of disciplines, including artificial intelligence, medicine and environmental science. Its rigorous peer-review process and high impact factor make it a trusted source for cutting-edge discoveries that shape scientific progress. The high citation rate of Nature articles reflects their far-reaching impact, often setting new research directions and informing policy and industry advances.
The presence of Journal of Medical Internet Research (JMIR) with 633 citations and The Cureus Journal of Medical Science with 448 citations highlights the academia interest in the use of AI, genAI and LLMs in healthcare
# Sources Articles
1 ARXIV 10,237
2 ADV NEUR IN 1,021
3 NATURE 980
4 J MED INTERNET RES 633
5 IEEE ACCESS 618
6 LECT NOTES COMPUT SC 464
7 COMPUT HUM BEHAV 454
8 CUREUS J MED SCIENCE 448
9 SCIENCE 448
10 INT J INFORM MANAGE 405
Table 2. Most Local Cited Sources
co-citation analysis, bibliographic coupling, keyword co-occurrence and other techniques
4.5. Most Relevant Countries
The Table 3 shows the country distribution of research articles based on corresponding authors, distinguishing between single-country publications (SCP) and multi-country publications (MCP). The USA leads with 610 articles (30.4%), of which 18% involve international collaboration (MCP). China (10.5%) follows, but with a higher proportion of MCPs (33.3%), indicating a stronger commitment to international research. The United Kingdom (7.4%) and Canada (2.9%) have some of the highest MCP percentages (45.9% and 41.4% respectively), reflecting their active participation in global collaborations. Other notable contributors include Germany (5.1%), Australia (4.2%) and India (3.2%), all of which also have a significant share of international partnerships, indicating the increasing globalisation of AI and LLM research. For better visualization, we also provide Figue 3.
This highlights the trend that the US and China are the leading forces driving innovation in AI and LLM research, reflecting their dominant roles in technological advancements and AI development, with technologies such as OpenAI’s ChatGPT in USA and DeepSeek in China. Meanwhile, the European Union (EU) tends to focus on regulation, as seen in policies such as the GDPR and the EU AI Act, emphasising ethical oversight and legal frameworks rather than direct competition in AI innovation.
# Country Articles Articles % SCP MCP MCP %
1 USA 610 30.4 500 110 18.0
2 China 210 10.5 140 70 33.3
3 United Kingdom 148 7.4 80 68 45.9
4 Germany 102 5.1 69 33 32.4
5 Australia 85 4.2 54 31 36.5
6 India 65 3.2 42 23 35.4
7 Canada 58 2.9 34 24 41.4
8 Spain 57 2.8 38 19 33.3
9 Italy 56 2.8 41 15 26.8
10 Netherlands 33 1.6 19 14 42.4
Table 3. Country Contribution Based on Corresponding Authors
Figure 3. Country Contribution Based on Corresponding Authors. Note: MCP represents multi-country publications, while SCP indicates the ratio of single-country publications.
4.6. Thematic Evolution
The thematic evolution of AI ethics and trustworthiness in LLM concerns is illustrated in the Figure 4, which maps the shifts in research focus from 2019 to 2025. Initially, between 2019 and 2022, research focused predominantly on broad terms such as artificial intelligence, ethics, machine learning, language models and AI – reflecting general concerns about AI development and ethical challenges.
However, in 2023-2024, the focus begins to narrow towards more specialised topics such as large language model, generative artificial intelligence, responsible AI and AI ethics. The emergence of these topics coincides with the surge in popularity of ChatGPT, but studies on the specific ethical or responsible dimensions of large language model and generative artificial intelligence are still in their infancy.
Regarding AI ethics and responsible AI, terms used in this period, they are discussed through theoretical principles (Jobin et al., 2019), (Floridi and Cowls, 2022), (Mittelstadt, 2019), (Vakkuri et al., 2021) and the urge to translate these principles into practice (Morley et al., 2019b), (Vakkuri et al., 2020), (de Cerqueira et al., 2022), (Siqueira De Cerqueira et al., 2021).
As AI systems – especially LLMs – became more advanced, concerns over trust, transparency, and accountability became even more pronounced. In 2025-2025, Figure 4 reveals a diversification of research topics, branching into AI in education, blockchain, trustworthiness, creativity, chatbots, and disinformation.
The rise of trustworthiness comes mainly from the term responsible AI, this demonstrates how the discourse has shifted from traditional responsible AI principles in AI ethics to what is trustworthiness in generative AI. This progression highlights the tension between innovation and ethical oversight, reinforcing the need for concrete trustworthiness frameworks rather than abstract trustworthiness taxonomies.
Similarly, the focus on AI ethics has shifted to concerns about disinformation, reflecting the fact that LLMs are rapidly replacing traditional search engines and are primarily used as chatbots, particularly in question-answering (QA) settings. As users increasingly rely on LLMs rather than traditional search engines to retrieve information, the risk of misinformation and manipulated content has become a central issue. This typical type of use of LLMs suggests why discussions of AI ethics have evolved to disinformation.
Figure 4 aligns with the bibliometric mapping approach of this study, revealing how AI ethics has been rebranded from abstract guiding principles into abstract trustworthiness taxonomies for LLMs.
Researchers that have never delved on Artificial Intelligence suddenly started researching about their ethical issues but in the generative AI ChatGPT era, forcing them to come up with new old terminologies that have been around for almost a decade. This sudden interest is, in the opinion of the first author of this paper, hindering the actual advancements towards more ethical AI, as it is masking the debate towards a ”new” set of theoretical guidelines as an enforced contribution, but not providing practical guidance for the developers, the ones actual responsible for the development of ethical AI.
In other words, this perceived rebranding of responsible AI and AI ethics to trustworthiness and disinformation is for the most part failing in proposing useful unified terminologies, frameworks, or practical guidance for developers to operationalize principles.
There is a growing need for enforceable guidelines beyond non binding ethical commitments. Even more evident after big tech companies such as Google withdrawing its AI ethical principles in mass surveillance and AI weapons (Hooker and Vallance, 2025). Consequently, there is an urge for actual hard laws that can govern the development and use of these systems, that can move beyond the EU AI Act (Sillberg et al., 2024), due to the fact that it only propose risk level assessments, and links to the AI HLEG (High-Level Expert Group on Artificial Intelligence, 2019). The later has a set of trustworthy AI principles, however, it is non binding. Despite the fact that a binding document links to a non binding one, does not automatically transform the AI HLEG into hard law. On top of that, an unified set of practical guidance are still lacking, after several years of debate – at least since 2019 (Morley et al., 2019b).
Despite this progress, fragmentation remains, with multiple authors competing with definitions of LLM trustworthiness rather than a single, unified framework.
The continued proliferation of non binding framework proposals and terminologies by singular researchers highlights the urgency of bridging the gap between theoretical AI ethics and its real-world implementation through unified frameworks and legislations.
Figure 4. Thematic Evolution: 2019-2025.
The word cloud in Figure 5 provides a visual representation of the most common keywords in the literature related to LLM, genAI and its trust and ethical implications. The size of each keyword indicates its frequency, with larger words appearing more frequently in the reviewed studies.
Among the most prominent terms, ”ChatGPT”, ”trust”, ”AI”, ”technology”, and ”information” are the most frequently mentioned, indicating that discussions around trust and AI adoption play a crucial role in ongoing research. It also shows that researchers are mainly making observations based on the ChatGPT platform, which is certainly the main LLM used in academia.
In addition, ”health” and ”education” emerge as recurring themes, reflecting concerns about the impact of AI in different domains, and broadening the debates about the responsible use of AI and its impact on different domains, particularly education and healthcare.
The word cloud also highlights other frequently cited themes, such as ”performance,” ”user acceptance,” ”privacy,” ”ethics,” ”impact,” ”challenges,” ”bias,” ”intention,” ”risk,” and ”opportunities.” These recurring themes highlight the growing emphasis on trustworthy AI, with concerns such as bias, privacy and ethics closely linked to user acceptance. At the same time, the literature recognises both challenges and opportunities, emphasising the need for responsible AI adoption while exploring its potential benefits.
Overall, Figure 5 illustrates the multifaceted nature of ChatGPT research, encompassing technical advances, ethical considerations, societal impacts, and real-world applications, particularly in areas such as trust in AI, automation, education, and user interaction. The findings reinforce the importance of ensuring ethical implementation of AI, transparent models and responsible innovation in this rapidly evolving field.
Figure 5. Word cloud.
4.7. Co-Occurrence Network by Keyword Plus
The co-occurrence network visualization in Figure 6 highlights the central role of ”artificial intelligence”, with strong associations to key themes like ”ChatGPT”, ”chatbot”, ”large language models”, related to the technologies themselves, and to ”medicine”, ”science”, ”education”, ”health”, relating to domains of interest; and ”ethics” and ”bias” emphisizing the ethical concerns. This suggests that AI research is deeply interconnected with societal and ethical challenges, particularly in LLMs and their impact on various domains.
On the other hand, the blue cluster focuses on the concept of ”trust”, linking it to ”acceptance”, ”user acceptance”, ”perceptions”, ”adoption”, ”quality” and ”privacy”. This indicates a growing interest in AI trustworthiness, highlighting it as in terms of user acceptance. That is, instead of having previous set of AI ethical guidelines governing AI use, development and deployment in a more holistic way, this suggests a shift in a more simplistic view of trust in AI is simply gauging the acceptance of a LLM by the user.
The co-occurrence network in Figure 6 highlights the the divide between ethical AI development and its trust/trustworthiness implications.
Figure 6. Co-Occurrence Network.
The clustering by coupling analysis presented in Figure 7 categorizes research themes based on author coupling, using references as a measure. The x-axis (Centrality) represents the extent to which a topic is interconnected with others, while the y-axis (Impact) reflects the influence of each cluster based on global citation scores. The clusters highlight key themes in AI research, with a strong focus on artificial intelligence, ChatGPT, trust, and performance. Notably, the red cluster (AI, trust) has high confidence values, suggesting a strong association between AI discussions and trustworthiness concerns.
The green cluster, positioned near the center, contains artificial intelligence, ChatGPT, and performance, indicating a balance between centrality and impact. This suggests that discussions about ChatGPT’s capabilities and AI performance are well-integrated within broader AI research. The top-right cluster, though smaller, highlights decision-making and AI systems, implying its emerging importance in AI governance. The analysis suggests a growing intersection between AI ethics, trust, and performance, reinforcing the need for responsible AI development and evaluation frameworks.
Our main interest resides in the red cluster, as it discusses trust. Table 4 shows the most important authors ranked by its relevance within the red cluster.
Overall, the five studies highlight the rapid adoption and diverse applications of large language models (LLMs) across multiple domains, while also highlighting their ethical, practical, and evaluative challenges. Martinez-Maldonado et al. systematically review the use of LLMs in education and identify nine key applications, such as grading, content generation and feedback, while cautioning about replicability, transparency and ethical concerns. Park et al. extend this discussion by investigating the impact of misinformation generated by ChatGPT on tourism decision-making, demonstrating how trust in AI can be easily manipulated when incorrect information is prominently presented. Similarly, Sallam examines AI in healthcare, highlighting its potential to improve medical research, education and clinical workflows, but also highlighting risks such as bias, hallucination, plagiarism and cybersecurity threats.
Wang et al. provide a meta-analysis of AI evaluation methods and argue for a more structured approach to assessing their performance, ethical considerations and real-world impact in different domains. This ties in with Herrera et al.’s work on Explainable AI (XAI), which highlights the urgent need for transparency in AI decision-making and proposes 28 open problems for achieving better interpretability and accountability in AI systems. Taken together, these studies show that while LLMs have transformative potential, their responsible use requires robust evaluation methods, ethical guidelines and accountability.
# Author Title Citation
1 Martinez-Maldonado R Practical and ethical challenges of large language models in education: A systematic scoping review (Yan et al., 2023)
2 Park J When ChatGPT Gives Incorrect Answers: The Impact of Inaccurate Information by Generative AI on Tourism Decision-Making (Kim et al., 2023)
3 Sallam M ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns (Sallam, 2023)
4 Wang C A Survey on Evaluation of Large Language Models (Chang et al., 2023)
5 Herrera F Explainable Artificial Intelligence (XAI) 2.0: A manifesto of open challenges and interdisciplinary research directions (Longo et al., 2024)
Table 4. Top Cited Authors within ’trust’ Cluster and Their Works
Figure 7. Clustering by Coupling.
4.8. The LLM trust/trustworthiness definitions proposed by different authors in the literature
In Table LABEL:tab:trust_definitions, we provide a structured overview of various definitions and conceptualizations of trust and trustworthiness in AI, as discussed in the literature. This compilation allows for a comparative analysis of how trust in AI is framed across different domains, emphasizing dimensions such as ability, reliability, integrity, explainability and ethical alignment. From the initial 68 analysed documents, 18 documents are explicitly or implicitly providing useful information to the task of framing what is trust/trustworthiness in LLM or genAI. That is, 26.47% against 73.53% of documents that are not providing useful information for this task. Other documents were not taken into consideration because they are either closed access, editorial, on the use of LLM to improve trust, in a different domain e.g., radiology, measuring trust level of a certain group when using certain LLM system, or simply with a different focus. Interesting to note that 5 of the discarded papers were related to fake news, misinformation and deepfake. The complete analysis is available on Zenodo.
Trust is a multidimensional construct that encompasses several key components, including benevolence, competence, and integrity (Mayer et al., 1995). Mayer et al.’s work entitled ”An Integrative Model of Organizational Trust” was published in 1995 in a Management venue, and has been adapted in several contexts, one of them being in trust in AI systems. We found a tendency in reviewed studies in utilizing Mayer et al.’s work to build on definitions for what is trustworthiness in LLM – 33,33%, one third, of definitions (6/18 entries: 5 (Wang et al., 2024a), 6 (Ganapati and Desouza, 2024), 8 (Lacity et al., 2024), 12 (Ressel et al., 2024), 15 (Jia et al.(2025), Jasper), 16 (Martell et al., 2024)) explicitly draw from Mayer et al.’s (Mayer et al., 1995) organizational trust framework. Its core elements – ABI+ factors of perceived trustworthiness – towards AI, are (Martell et al., 2024):
  1. (1)
    Ability: AI’s capacity to perform tasks effectively,
  2. (2)
    Benevolence: AI’s intent to act in users’ best interests,
  3. (3)
    Integrity: adherence to ethical standards,
  4. (4)
    Predictability: consistency in AI responses.
Trust in AI, particularly in situations of uncertainty and vulnerability, is described as the belief that an AI agent will support an individual in achieving his or her goals (Wang et al., 2024a). Trust is fundamentally linked to vulnerability and uncertainty, as highlighted by multiple perspectives. Hoy and Tschannen-Moran (Hoy and Tschannen-Moran, 1999) define trust as a willingness to be vulnerable based on confidence in another party’s benevolence, reliability, competence, honesty, and openness, a concept applied to AI-mediated teacher-student relationships. Mayer et al. (Mayer et al., 1995) reinforce this by framing trust as a positive expectation of another’s intentions and behavior, emphasizing the risk involved in relying on an entity. Similarly, Lee and See (Lee and See, 2004) describe trust as the belief that an agent will assist in achieving one’s goals despite uncertainty and vulnerability, underscoring that trust in AI systems depends not just on their performance but also on users’ confidence in their intent and reliability.
In this context, users assume the vulnerable position of relying on genAI models or proprietary platforms with the expectation that this risk will be beneficial to them, i.e. that it will help them achieve their goals despite possible misinformation or bias (Ressel et al., 2024). In the paper ”Mitigative Strategies for Recovering From Large Language Model Trust Violations” (Martell et al., 2024), the authors present three different levels of trust calibration, as shown in Table 5.
Table 5. Types of Trust Calibration in Human-AI Interaction (Martell et al., 2024)
Concept and Definition Reference
Appropriate (Calibrated) Trust: Alignment between an individual’s perceived trust level and the AI system’s actual capabilities and performance. (Yang et al., 2020)
Over-Trust: Occurs when users: Perceive the AI as more capable than it truly is Treat the technology as a fully competent teammate Over-rely on AI outputs despite known limitations (Aroyo et al., 2021)
Under-Trust: Occurs when users: Underestimate the AI system’s capabilities Over-rely on their own judgment Dismiss potentially valuable AI contributions (Hoff and Bashir, 2014)
The same study highlights the need to prioritise accuracy in LLM deployment, as early errors can irrevocably damage user confidence, calibration and subsequent adoption. In other words, operationalising trustworthiness in AI-LLM systems from the early stages is fundamental, as it is difficult to regain user trust. This can be seen in cases such as Microsoft’s Tay Bot  [^2] https://blogs.microsoft.com/blog/2016/03/25/learning-tays-introduction/. This AI-powered chatbot started tweeting offensive neo-Nazi comments within hours of its release and was subsequently shut down.
The more familiar we are with a tool, the more we tend to trust them (Califano et al., 2024). Other authors state that trust in AI-LLM can be translated as the degree to which users and institutions are confident that the data used by, and the information generated by, AI-LLM is impartial, i.e. unbiased (Jenks, 2024).
Two more trust related dimensions are presented in Table LABEL:tab:trust_definitions: cognitive trust (rational evaluation of a system’s reliability and accuracy) and affective trust (emotional confidence in its benevolence and alignment with user interests) (Zhang et al., 2014). Cognitive trust, anchored in perceived credibility, facilitates affective trust through sustained reliance, as users transfer trust from the platform’s technical competence to its perceived ethical intent (Zhang et al., 2014; Wang et al., 2016). This interdependence highlights how platform reliability shapes users’ perceptions of AI-generated content, with technical accuracy (’reliable’) and empathetic design (’credible’) jointly calibrating holistic trust.
What are the trade-offs of using or adapting a 30 year old organizational trust framework to the AI era? First, as a benefit, it is a well established set of trust taxonomy, that has been widely accepted in the academia. However, it was devised way before wider adoption of different AI technologies, in special LLMs. Similarly to AI ethics, where soft law recommendations did not pose serious changes, should trustworthiness in LLM discussion be on hard laws that can unify the terms and recommendations?
In (Jenks, 2024), the authors state three principles of AI-LLM: fair, unbiased, and ethical. On top of that, they relate these principles with the work done within the European Commission on AI ”Building Trust in Human-Centric Artificial Intelligence” (Commission, 2019), which states that trustworthy AI should not only be lawful, ethical, and robust, but also give humans the ability to have control over its use, draw from and provide reliable data, offer security and privacy to its users, and be transparent, accountable, and inclusive. The efforts of the European Commision on AI is better translated in the EU AI Act (Sillberg et al., 2024; Commission, 2023), where a risk level assessment for AI systems is provided. We briefly describe the EU AI Act’s risk level in Table 6. From this Table, we pinpoint that most of genAI and LLM generated content falls under the ”Limited Risk” level, where there is a need for Transparency requirements, highlighting once again the importance of this trustworthy principle.
Risk Level Description
Unacceptable Risk AI systems that are prohibited due to threats to fundamental rights (e.g., social scoring, manipulation).
High Risk AI systems that impact safety or fundamental rights (e.g., biometric ID, credit scoring, law enforcement).
Limited Risk AI systems with transparency requirements (e.g., chatbots, AI-generated content).
Minimal Risk AI systems with no specific regulations (e.g., spam filters, AI-driven recommendations).
Table 6. Risk Levels in the EU AI Act
In addition, within the EU AI Act there is reference to a non-binding document, also provided by the European Comission, that aims to provide a set of trustworthy principles in AI. This non-binding document is the 2019 Ethics Guidelines for Trustworthy AI developed by the AI High-Level Expert Group (HLEG) (High-Level Expert Group on Artificial Intelligence, 2019). We shortly describe its principles in Table 7. This Table should complement our Table 9, as there are many overlapping terms, such as Transparency (Transparency) and Accountability (Responsible).
Principle Description
Human Agency and Oversight AI should respect human autonomy and be controllable.
Technical Robustness and Safety AI should be resilient, secure, and function as intended.
Privacy and Data Governance AI must ensure data protection and quality.
Transparency AI should be explainable and provide clear information.
Diversity, Non-discrimination, and Fairness AI should be inclusive and avoid bias.
Societal and Environmental Well-being AI should promote sustainability and benefit society.
Accountability AI should have mechanisms for responsibility and oversight.
Table 7. AI HLEG Principles for Trustworthy AI
Finally, in (Baeza-Yates, 2024), there is a critic towards the use of the terms trustworthy and ethical AI, as they can be used to humanize, i.e., anthropomorphise AI. This critic is valid insofar every Information System artefact, such as LLM and AI systems, are devised by humans, and they are programmatically designed to fulfil a set of subjective or objective desires implied by the developers or organizations they are inserted.
Table 8. Trustworthiness Definitions in AI
# Trust/trustworthiness definition Synonyms to trustworthy AI Ref Builds on
1 Foundation Models (FMs) and Large Language Models (LLMs) demonstrate strong generalization capabilities and adaptability to specific tasks through fine-tuning or in-context learning. However, they currently lack the predictability and robustness required for deployment ’as is’ in adversarial environments. In this context, trustworthiness is characterized by the ability to be directly deployed in real-world scenarios while maintaining resilience against adversarial conditions. Trustworthy, resilient, and interpretable artificial intelligence (AI) (Jha et al., 2023) -
2 To improve trust (in these models) is to understand the behavior of deep neural networks for Information Retrieval (IR). Explainable AI (Lucchese et al., 2023) -
3 Not explicitly, but states that: LLM’s ability in providing reliable (S&P) advice. Trusted, Reliable (Chen et al., 2023) -
4 The authors argue that the perceived trustworthiness of GAN-generated content depends on several factors, including the realism of the content, the context in which it is presented, and the pre-existing trust individuals have in the sources of information. More specifically in the case of fake news and misinformation. The authors suggest that addressing the ethical and political dimensions of GANs is crucial for maintaining trust in democratic societies. Perceived trustworthiness, disinformation concerns (Carnevale et al., 2023) -
5 Trust in AI is described as the belief that an AI agent will support an individual in achieving their goals, particularly in situations marked by uncertainty and vulnerability. Trust plays a crucial role in user interaction with AI, as its absence can lead to reluctance in adopting AI tools, even when they demonstrate superior performance. Conversely, excessive trust in AI, especially in high-stakes domains like software engineering, can cause users to overlook potential errors or risks. The study explores strategies for fostering appropriate trust in generative AI tools, such as GitHub Copilot, to enhance their usability and reliability. Ability, Benevolence, Integrity (Wang et al., 2024a) (Lee and See, 2004) (Liao and Sundar, 2022) (Vereschak et al., 2021) (Commission, 2019) (Das and Rad, 2020) (Boubin et al., 2017) (O’Connor et al., 2019) (Pearce et al., 2022) (Perry et al., 2023) (Mayer et al., 1995)
6 Trust in AI comprises three key components: ability (ensuring reliable performance), integrity (aligning with ethical and social norms), and benevolence (acting in society’s best interest). Trustworthiness is reinforced by verifying AI’s reliability, ensuring ethical alignment, and strengthening oversight through public scrutiny and regulation, emphasizing governance over technical solutions. Ability, integrity, benevolence, verifiability (Ganapati and Desouza, 2024) (Desouza and Dawson, 2023) (Mayer et al., 1995)
7 While not formally defining trustworthiness, it conceptualizes it as the ability of AI systems to avoid generating false or misleading responses by accurately determining when they lack sufficient evidence to answer a given question. Reliable, factual (Łajewska and Balog, 2024) -
8 Trust in IT artifacts is defined by their quality, reliability, accuracy, and security. While organizational trust frameworks, such as Mayer et al. (Mayer et al., 1995), are widely used in IS research, applying concepts like benevolence to IT artifacts risks anthropomorphizing technology. Ability/competence, integrity, and benevolence (Lacity et al., 2024) (Mayer et al., 1995) (Porra et al., 2019) (Belanger and Crossler, 2019) (Kim et al., 2009) (Lowry et al., 2013)
9 Not explicitly, but it describes trustworthy AI in terms of consistency, reliability, explainability, and safety. Consistency, reliability, explainability, and safety (Gaur and Sheth, 2024) -
10 It follows a definition of trust in education, which states that: ”Trust is an individual’s or group’s willingness to be vulnerable to another party based on the confidence that the latter party is benevolent, reliable, competent, honest, and open.” This definition is applied to teacher-student relationships to explore how trust is built and eroded in an AI-mediated assessment environment. Transparency, reliability, honesty, competence, integrity, fairness, benevolence, openness, trust-building (Luo, Jess) (Hoy and Tschannen-Moran, 1999)
11 Believing in the credibility or accuracy of a statement. Credibility, AI alignment (Buchanan and Hickman, 2024) -
12 Trust is characterized by a willingness to be vulnerable, accompanied by a positive expectation of the other party’s intentions and behavior. In the context of AI and LLMs, real-world incidents have urged the need for translating high-level trustworthiness rhetoric into domain-specific practices. This raises the question of how trust should be defined in fields such as the insurance industry and whether regulatory bodies should actively promote trust in AI systems. AI ethics, AI trustworthy (Ressel et al., 2024) (Mayer et al., 1995)
13 While trust in AI-generated content is not explicitly defined, the authors suggest that trust can be enhanced by providing indicators of familiarity, validation, or endorsements from experts. However, they also note that people tend to be more cautious when AI-generated content is explicitly labeled. Transparency, disclosure, validation, familiarity, experience, task suitability, credibility (Califano et al., 2024) -
14 Trust in AI-LLM is defined as confidence in the impartiality of its data and generated information. Personalization plays a crucial role in fostering trust in AI-LLM, particularly in intercultural communication. According to the European Commission (2024), trustworthy AI should not only be lawful, ethical, and robust but also empower users with control over its use, ensure data reliability, security, and privacy, and uphold transparency, accountability, and inclusivity. Ultimately, trust in AI-LLM depends on the perception that its outputs are reliable and unbiased, with trustworthiness requiring ethical, inclusive, and accountable AI-generated knowledge. Confidence, fair, unbiased, ethical, lawful, robust, human oversight, transparency, accountability, security, privacy (Jenks, 2024) (Cook, 2001) (Araujo et al., 2020) (Commission, 2025)
15 Trust is a multidimensional construct that encompasses several key components, including benevolence, competence, and integrity. Research suggests that users tend to trust AI-generated content less when they are aware of its artificial origin, a trend observed in multiple studies. Credibility, Transparency (Jia et al.(2025), Jasper) (Mayer et al., 1995)
16 Trust is defined as the belief that an agent will assist in achieving an individual’s goals in situations marked by uncertainty and vulnerability. Appropriate trust, or calibrated trust, occurs when a person’s trust in a system aligns with its actual performance. Over-trust arises when individuals overestimate AI capabilities, while under-trust occurs when users undervalue AI and rely too heavily on themselves. The ABI+ framework outlines four key dimensions of trust: (1) Ability – AI’s capacity to perform tasks effectively, (2) Benevolence – AI’s intent to act in users’ best interests, (3) Integrity – adherence to ethical standards, and (4) Predictability – consistency in AI responses. This highlights the importance of ensuring AI systems maintain accuracy from early deployment stages, as initial errors can have long-term consequences on user trust calibration and later adoption. Transparency, calibrated trust, reliability (Martell et al., 2024) (Mayer et al., 1995) (Lee and See, 2004) (Yang et al., 2020) (Aroyo et al., 2021) (Hoff and Bashir, 2014) (Dietz and Den Hartog, 2006) (Toreini et al., 2020)
17 Trust in generative AI can be divided into cognitive trust and affective trust. Cognitive trust is based on a user’s evaluation of the AI platform’s ability to deliver accurate and reliable information, while affective trust stems from the user’s belief that the AI system considers their interests and well-being. Reliable, credible (Zhou and Lu, 2024) (Zhang et al., 2014) (Wang et al., 2016) (Stewart, 2003)
18 The author actually criticizes the term trustworthy and ethical AI because they can humanize AI. Responsible AI (Baeza-Yates, 2024) -
In Table 9, we present the frequency of key terms associated with trustworthiness in AI, reflecting the most commonly emphasized aspects in the literature. Transparency and Explainability appear as the first and second most frequently mentioned characteristics, respectively, highlighting how they are enabling principles, i.e., without transparency and explainability it is impossible for other trustworthiness or ethical principles to exist (Siqueira De Cerqueira et al., 2021). Reliability is related to Ability and Predicability in the ABI+ framework (Mayer et al., 1995). Furthermore, Ethics, Privacy, and Fairness highlight concerns related to Integrity, i.e., adherence to ethical standards. This Table 9 provides a quantitative perspective on the many trustworthiness aspects identified from the literature.
Term Frequency
Transparency 12
Explainability 9
Reliability 9
Ethics 6
Privacy 6
Fairness 5
Responsible 5
Credibility 4
Robustness 4
Trustworthy 4
Table 9. Frequency of Trustworthiness-Related Terms
4.9. Practical LLM trustworthiness-enhancing techniques found in the literature
From the 68 documents analysed, 19 unique documents contribute to the creation of the Table LABEL:tab:trust_enhancing, which provides strategies for enhancing trustworthiness in LLM. We found 20 different strategies, categorised by their LLM lifecycle phase – either pre-training, post-training or inference phase – the actor responsible for their implementation – either developer or user – and the trustworthiness attributes associated with them – e.g. transparency, explainability, reliability.
In the first place, with 6 different documents referring to this strategy, is (supervised) fine-tuning (SFT), in the post-training phase. This strategy is the responsibility of the LLM system developer and focuses mainly on reliability, robustness and fairness. This is justified by the boom of chatGPT, which is a fine-tuned version of the GPT-3 or GPT-4 pre-trained models. SFT has enabled models to become helpful assistants, that is, only the pre-trained models themselves were not the desired chatbot that chatGPT came to be. Nowadays, there is extensive documentation such as OpenAI fine-tuning docs, and platforms such as Google Colab, which have facilitated the availability of hardware and environment to allow fine-tuning by anyone anywhere in the world.
In second place, with 4 different documents referring to this strategy, is Retrieval Augmented Generation (RAG), which also takes place in the post-training phase and under the responsibility of the developer. This technique has proved to be efficient, cost-effective and allows models to base their responses on frequently updated documents. This strategy has demonstrated versatility and requires fewer technical resources to implement. The trustworthiness attributes most associated with RAG are faithfulness, transparency and reliability.
Developers have primary responsibility for trustworthiness (16/20 techniques), while users contribute minimally (4/20, e.g., human-in-the-loop feedback). This raises concerns about scalability, as techniques such as stakeholder feedback (Entry 5) require sustained user engagement, which is rarely seen in real-world deployments.
Users can improve the trustworthiness of AI-LLM systems mainly in the inference phase, through human-in-the-loop (evaluating the output), stakeholder feedback (expert feedback), prompt engineering (implementing prompt techniques) and Reinforcement Learning from Human Feedback (RLHF) (providing feedback/labelling to the generated output).
Most techniques (12/20) target the post-training phase (e.g. fine-tuning, RAG, XAI methods). This overemphasis on post-training strategies rather than pre-training (e.g., bias mitigation through pre-training data curation) should not always suggest that developers are attempting to address reliability risks only after deployment. This is due to the fact that the pre-training phase (e.g., data gathering, training) typically requires millions of dollars and supercomputers. However, it is suggested from previous RQs that trust issues should be addressed before deployment, as early failures can severely damage user confidence and adoption, making it difficult for a platform to regain its image and user trust (Martell et al., 2024). In this study, we suggest that the model has passed the deployment phase only when it is in the inference phase.
Explainability and transparency dominate (14/20 entries), driven by techniques such as contextual explanations (Entry 8) and uncertainty expressions (Entry 11). In contrast, security (Entry 16: blockchain) and ethics (Entry 13: audits) receive limited attention, indicating a misalignment with regulatory priorities such as the EU AI Act’s emphasis on systemic risk management.
The strategies identified suggest the need for integrated frameworks that unify technical interventions (e.g. RAG, knowledge graphs) with governance mechanisms (e.g. post-release audits, human-in-the-loop) across the LLM lifecycle, with a strong adherence to legislations.
Efforts such as the EU AI Act From the policy making perspective, it is pressing to enact regulations to mandate transparency and accountability in the development and deployment of LLMs.
Table 10. Trustworthiness-Enhancing Strategies in LLMs
# Trustworthiness-Enhancing Strategy LLM Lifecycle Phase Responsibility Trustworthiness Attributes Citations
1 (Supervised) Fine-tuning Post-training Developer Reliability, Robustness, Fairness (Jha et al., 2023) (Kilhoffer and Bashir, 2024) (Barman et al., 2024) (Ressel et al., 2024) (Chen and Shu, 2024) (Roumeliotis et al., 2025)
2 RAG Post-training Developer Faithfulness, Transparency, Reliability (Jha et al., 2023) (Lan et al., 2024) (Ressel et al., 2024) (Hannah et al., 2025)
3 Training Pre-training Developer Robustness, Accuracy (Bhattacharya et al., 2023)
4 Human-in-the-loop Inference User Fairness, Transparency, User Trust (Wang et al., 2024b) (Chen and Shu, 2024)
5 Stakeholder feedback Inference User Trust Calibration, Social Acceptability (Wang et al., 2024b)
6 Prompt engineering Inference User Safety, Controllability, Fairness (Mousavi et al., 2024) (Hannah et al., 2025)
7 Evaluation Post-training Developer Accountability, Transparency (Mousavi et al., 2024) (Nag et al., 2023)
8 Contextual Explanation Pre-training/Post-training Developer Explainability, Transparency (Chari et al., 2023)
9 In-context Learning (ICL) Post-training Developer Adaptability, Efficiency, Fairness (Zhang et al., 2024) (Barman et al., 2024)
10 RAG-Ex: a model and language-agnostic explanation framework Post-training Developer Explainability, Transparency (Sudhi et al., 2024)
11 Uncertainty Expression in LLMs Post-training Developer Transparency, Explainability, Reliability (Kim, 2024)
12 LLMs’ self-explanations (i.e., explanations of their own answers and behaviors) Post-training Developer Explainability, Interpretability (Kim, 2024)
13 Post-release audit tool as an evaluative framework Post-training Developer Transparency, Accountability, Explainability, Ethics (Fernandez Nieto, 2024)
14 Reinforcement Learning from Human Feedback (RLHF) Inference Developer/User Alignment (Barman et al., 2024) (Chen and Shu, 2024)
15 Chain of Thought (CoT) Inference Developer Explainability (Barman et al., 2024) (Chen and Shu, 2024)
16 Blockchain Post-training Developer Security, Transparency, Tamper-resistance (Fan et al., 2024)
17 Knowledge Graph Post-training Developer Transparency, Explainability (Chen and Shu, 2024) (Hannah et al., 2025)
18 (Training) Data curation Pre-training Developer Fairness, Bias Mitigation (Chen and Shu, 2024)
19 (Access to external) Tools (e.g., access to search engines online) Inference Developer Explainability (Chen and Shu, 2024)
20 eXplainable AI (XAI) techniques – model simplification (LIME), perturbation-based methods (SHAP), gradient-based approaches (InputXGradient, Grad-CAM), Layer-wise Relevance Propagation (LRP) Post-training Developer Explainability, Interpretability, Transparency (Mersha et al., 2025)
5. Limitations
While bibliometric analysis provides valuable insights into research trends, it has notable limitations. Its emphasis on publication and citation counts does not always capture the depth, quality or maturity of a field of research (Donthu et al., 2021). High citation rates can indicate popularity rather than true intellectual impact, and biases in citation patterns can distort impact (Donthu et al., 2021).
To address these concerns, this study combines bibliometric techniques with qualitative assessments of articles with high focus on discussing trustworthiness (i.e., studies with ”trust*” in the abstract). By integrating both perspectives, we aim to provide a more comprehensive understanding of LLM trustworthiness research, including also more detailed, qualitative analysis of select papers in addition to the quantitative bibliometric approach.
Another limitation of this study is that it relies solely on data from the Web of Science database. This results in two limitations. First, given the rapid pace of research in LLMs, a significant proportion of relevant studies are often published on arXiv, a widely used open-access preprint repository maintained by Cornell University. By not including arXiv, this study may not capture the latest developments and trends in the field. However, by not including them, this study aims to focus only on peer-review papers. Second, by only focusing on one database, we will also have excluded some potentially relevant, peer-reviewed and already published research. However, this was done due to technical limitations related with the chosen bibliometric approach.
6. Final Remarks
This study provides a bibliometric analysis of trustworthiness in Large Language Models (LLMs), highlighting key research trends, influential papers and gaps in the literature. The findings reveal a shift from traditional AI ethics discussions towards trustworthiness frameworks, but a lack of consensus and practical implementation strategies persists. While various approaches such as transparency, accountability and governance mechanisms have been proposed, enforceable regulations and standardised frameworks remain necessary for real-world impact.
Future research should focus on bridging the gap between theoretical taxonomies and actionable trustworthiness-enhancing techniques to ensure their practical application in LLM development and deployment. In addition, interdisciplinary collaboration is essential to reconcile ethical principles with technical and regulatory considerations. Given the rapid evolution of LLMs and their increasing societal impact, ongoing efforts must prioritise both responsible innovation and policy-driven accountability. Ultimately, achieving trustworthy AI will require not only technological advances, but also robust governance, public engagement, and industry-wide commitment to ethical AI development.
Acknowledgements.
This research was supported by Jane and Aatos Erkko Foundation through CONVERGENCE of Humans and Machines Project under grant No. 220025.